Corporate finance

nFlashback Corporation is evaluating an extra dividend versus a share repurchase. In either case, $12000 would be spent. Current earnings are $4 per share, and the stock currently sells for $100 per share. There are 4000 shares outstanding. No tax. 

na) evaluate the two alternatives in terms of the effect on the price per share and shareholder wealth

nwhat will be the effect on Flashback's EPS and PE ratio under two different scenarios?
Answer #1

Given in the question:


Value of Extra Dividend = Value of Share Repurchase = $ 12000


Current EPS = 4


Current Share Price, P = 100


No. of Shares Outstanding, N = 4000


We can calculate Earnings from EPS and the no. of shares outstanding;


EPS = Earnings / No. of Shares


Therefore,


Earnings = EPS * No. of Shares Outstanding = 4*4000 = 16000


Case I: Extra Dividends


If $ 12000 is spent on giving dividends and the no. of shares outstanding are 4000, then;


Dividend per Share = 12000/4000 = 3.00 ............................ (Pt. No. 1)


(a) Effect on Price of Share


In case of cash dividend, the New Share Price per share will decrease by the amount of DIvidend per share.


So,


New Share Price = Old Share Price - Dividend Per Share = 100 - 3 = $ 97


(b) Effect on EPS


After the declaration of dividends, earnings will remain the same, and also the no. shares outstanding will remain the same. So, there will be no effect on the EPS after a cash dividend.


(c) Effect on Shareholder Wealth


Earlier, that is before the payment of Cash Dividend, a shareholder owning 1 share had wealth equal to the share price;


Old Wealth = Share Price = 100


After dividend payment, the same shareholder will now have wealth equal to the New Share Price + Cash Dividend per share;


New Wealth = 97 + 3 = 100


So, there is no effect on shareholder wealth due to the payment of cash dividends.


(d) Effect on P/E Ratio


Old P/E Ratio = Share Price/EPS = 100/4 = 25


We have seen in part (b) that the Earnings per share will not be affected due to the payment of cash dividends.


So, New P/E Ratio = New Share Price/Earnings Per Share = 97/4 = 24.25


So, P/E Ratio decreases due to the payment of cash dividends.


Case II: Share Repurchase


If $ 12000 is spent on share repurchase and the share price is $ 100 per share, then the no. of shares repurchased = 12000/100 = 120, therefore now the New no. of shares outstanding = (4000 - 120) = 3880,


(a) Effect on Price of Share


Share Price before Share Repurchase = 100


New Share Price = Market Value of Equity after Repurchase/Outstanding no. of Shares after Repurchase


Market Value of Equity after Repurchase = (No. of Shares Outstanding earlier * Share Price Earlier) - Amount of Repurchase


= (4000 * 100) - 12000


= 388000


So,


New Share Price = 388000/3880 = 100 $


So, there is no effect on Share Price due to the repurchase.  


(b) Effect on EPS


After the repurchase, earnings will remain the same and the no. shares outstanding will decrease. New SHares Outstanding = 3880. ................ (Has been calculated earlier)


New EPS = Earnings / New No. of Shares = 16000/3880 = 4.12.


So, The EPS has increased from 4 to 4.12 due to share repurchase.


(c) Effect on Shareholder Wealth


Before repurchase, a shareholder owning 1 share had wealth equal to the share price;


Old Wealth = Old Share Price = 100


After repurchase, the shareholder wealth again will be equal to the New Share Price;


New Wealth = New Share Price = 100


So, there is no effect on shareholder wealth due to share repurchase.


(d) Effect on P/E Ratio


Old P/E Ratio = Share Price/EPS = 100/4 = 25


We have seen in part (b) that the Earnings per share have increased from 4 to 4.12 due to the share repurchase.


So, New P/E Ratio = New Share Price/New Earnings Per Share = 100/4.12 = 24.27


So, P/E Ratio decreases due to sharing repurchase.


answered by: Cydney